# Possible molecules of triple-heavy pentaquarks within the extended local hidden gauge formalism

### v1

@@ Abstract
In the present work, we investigate the meson-baryon interactions of the open heavy sectors to search for the possible triple-heavy molecular pentaquarks, where the meson-baryon interaction potentials are obtained with the vector meson exchange mechanism within the extended local hidden gauge formalism. The scattering amplitudes are evaluated by solving the coupled channel Bethe-Salpeter equation, where some bound systems are found. With the poles of the amplitudes in the complex plane, the masses and widths of these bound states are obtained. Furthermore, we also calculate the couplings and compositenesses of different channels in each bound system to check their molecular nature. We predict four $\Omega_{ccc}$-like states, four $\Omega_{bbb}$-like states, fourteen $\Omega_{bcc}$-like states, and ten $\Omega_{bbc}$-like states, which can be hunted experimentally in future.
@@ I Introduction
Over the last twenty years, with the progress of high-energy physics experiments, a large number of new hadronic states have been discovered. It is found that many of them cannot be treated as the conventional hadronic states of mesons, which consist of a quark and an antiquark, or of baryons, which make up three quarks (antiquarks) Gell-Mann:1964ewy ; Zweig:1981pd . These states could be the candidates of the exotic states, such as glueballs, hybrids, multiquark states, etc., which catch much attention both in experiment and theory Liu:2013waa ; Hosaka:2016pey ; Chen:2016qju ; Richard:2016eis ; Lebed:2016hpi ; Olsen:2017bmm ; Guo:2017jvc ; Liu:2019zoy ; Brambilla:2019esw ; Meng:2022ozq ; Chen:2022asf . It is a key issue to deepen our understanding of the non-perturbative behavior of the strong interaction.
Among those reported states, a major group of the hidden-charm hadrons, such as the charmonium-like $XYZ$ states and the $P_{c}/P_{cs}$ pentaquark-like states, have stimulated an extensive discussion on the molecular multiquark states. As the first experimentally observed charmonium-like $XYZ$ state, the $X(3872)$ was announced by the Belle collaboration in the $B^{\pm}\rightarrow K^{\pm}\pi^{+}\pi^{-}J/\psi$ decays in 2003 Belle:2003nnu . Since its mass is so close to the $D\bar{D}^{*}$ threshold, it is naturally explained as the $D\bar{D}^{*}$ molecular state in Refs. Swanson:2003tb ; Wong:2003xk ; Close:2003sg ; Voloshin:2003nt ; AlFiky:2005jd ; Liu:2008fh ; Thomas:2008ja ; Liu:2009qhy ; Lee:2009hy ; Gamermann:2009uq ; Braaten:2010mg ; Baru:2013rta ; Baru:2015nea ; Song:2023pdq . In 2013, a charge charmonium-like state $Z_{c}(3900)$ was reported by the BESIII and Belle collaborations BESIII:2013ris ; Belle:2013yex , which is also close to the $D\bar{D}^{*}$ threshold and can be explained as a molecular state Guo:2013sya ; Wang:2013daa ; Aceti:2014uea . In 2015, the states $P_{c}(4380)^{+}$ and $P_{c}(4450)^{+}$ were discovered in the $J/\psi p$ invariant mass spectrum of the $\Lambda_{b}^{0}\rightarrow J/\psi pK^{-}$ decay by the LHCb collaboration LHCb:2015yax ; LHCb:2016lve , which were first predicted as the molecular pentaquark states in Refs. Wu:2010jy ; Wu:2010vk ; Wang:2011rga ; Yang:2011wz ; Wu:2012md ; Xiao:2013yca ; Li:2014gra ; Karliner:2015ina and attracted much theoretical attentions for the molecular states, see the discussions in the reviews Hosaka:2016pey ; Chen:2016qju ; Olsen:2017bmm ; Guo:2017jvc and references therein. Furthermore, in 2019, a fine $J/\psi p$ mass spectrum was observed by the LHCb collaboration LHCb:2019kea , where the two $P_{c}$ states found before were in fact three ones, the $P_{c}(4312)^{+}$, $P_{c}(4440)^{+}$ and $P_{c}(4457)^{+}$, and which provided strong support for the molecular picture Chen:2019asm ; Chen:2019bip ; Liu:2019tjn ; He:2019ify ; Xiao:2019mst ; Guo:2019kdc ; Xiao:2019aya . In 2021, the LHCb collaboration reported a hidden-charm pentaquark structure with strangeness $P_{cs}(4459)^{0}$ in the $\Xi_{b}^{-}\rightarrow J/\psi\Lambda K^{-}$ decay process LHCb:2020jpq . Moreover, in Ref. LHCb:2022ogu , a new state $P_{cs}(4338)^{0}$ containing a strange quark was observed in the decay $B^{-}\rightarrow J/\psi\Lambda\bar{p}$. These two states were considered most likely to be the $\bar{D}^{*}\Xi_{c}$ and $\bar{D}\Xi_{c}$ molecular states, respectively Xiao:2021rgp ; Du:2021bgb ; Zhu:2021lhd ; Lu:2021irg ; Zou:2021sha ; Wang:2021itn ; Wu:2021caw ; Yan:2022wuz ; Meng:2022wgl ; Burns:2022uha ; Ortega:2022uyu . Thus, under the molecular picture with the flavor symmetry, it is natural that the $P_{c}(4312)$ and $P_{c}(4440)$, $P_{c}(4457)$, are assigned to be the $\bar{D}\Sigma_{c}$ and $\bar{D}^{*}\Sigma_{c}$ states, respectively, and the $P_{cs}(4338)$, $P_{cs}(4459)$ to be the $\bar{D}\Xi_{c}$, $\bar{D}^{*}\Xi_{c}$ molecules, respectively, where they are just with the replacement of the light quark by the strangeness quark, $q\to s$.
@@
Note that, from the molecular picture, the hidden-charm pentaquark states $P_{c}$ and $P_{cs}$ can be regarded as a simple and natural substitution of $\bar{q}\rightarrow qq$ from the $XYZ$ states, as shown in Fig. 1. Analogously, under another natural substitution of $\bar{c}\rightarrow cc$ in the $XYZ$ states, the triple-charm pentaquark states can be obtained from the interaction of the charmed meson and the double-charm baryon, which catches our attention. In Refs. Chen:2017jjn ; Wang:2019aoc , the authors addressed this issue and predicted some isoscalar triple-charm molecular-type pentaquark candidates $D^{(*)}\Xi_{cc}$, $D_{1}\Xi_{cc}$, and $D_{2}^{*}\Xi_{cc}$ via the one-boson-exchange model, which can provide valuable information for future experiments. In view of their predictions and the motivation of the findings of the $P_{c}$ and $P_{cs}$ states, we make a further investigation of the triple-heavy systems in the present work, such as the pentaquark systems with the quark contents $cccq\bar{q}$, $bbbq\bar{q}$, $bccq\bar{q}$, and $bbcq\bar{q}$, which are referred as the $\Omega_{ccc}$-like, $\Omega_{bbb}$-like, $\Omega_{bcc}$-like, and $\Omega_{bbc}$-like molecular states. For the meson-baryon interaction, we adopt the coupled channel approach with the transition potential from the extended local hidden gauge formalism. Using such formalism, many molecular pentaquark states with double heavy quarks were predicted in Refs. Wang:2022aga ; Wang:2023mdj ; Song:2024yli .
Our paper is constructed as follows. In Sec. II, we introduce the $S$-wave interactions of the meson-baryon derived from the extended local hidden gauge formalism. Then, in Sec. III, we give the results of the triple-heavy pentaquark candidates by solving the the coupled channel Bethe-Salpeter equation. Finally, a short summary is shown in Sec. IV.
@@ II Formalism
Note that we only use the notation of the interaction meson fields for the vertex information in the vector exchange processes of Fig. 2, not the SU(5) symmetry. Therefore, the vertices $VPP$ and $VVV$ can be obtained from the effective Lagrangians of Eqs. (1) and (2).
However, for the vertex of $VBB$, it is difficult to directly extend the corresponding baryon fields matrix from the flavor SU(3) notation to the one of SU(5) with the charm and beauty baryon fields. Therefore, instead of the evaluation from the effective Lagrangian, we use the scheme of Ref. Debastiani:2017ewu to calculate the $VBB$ vertex from the flavor and spin wave functions of baryons and vector mesons, of which the results are consistent with the way from the effective Lagrangian. The baryons involved in our work, along with their explicit flavor and spin wave functions, are listed in Table 2, where $\chi_{MS}$ indicates that the spin wave function is mixed symmetric, and $\chi_{MA}$ represents the mixed antisymmetric spin wave function. For the $J^{P}=\frac{1}{2}^{+}$ ground baryon states in the case of $S_{z}=+\frac{1}{2}$, the corresponding spin wave functions are given by
@@
Based on above the local hidden gauge Lagrangians for the $VPP$ and $VVV$ vertices, and the extended approach for the $VBB$ vertex with the baryon wave functions, we obtain the general form of the tree level potentials after projecting to the $S$ wave
@@
Not that the coefficients of $C_{ij}$ in Table 3 are under the isospin basises, where we use the phase convention $|D^{(*)+}\rangle=|1/2,1/2\rangle$, $|D^{(*)0}\rangle=-|1/2,-1/2\rangle$, $|\bar{B}^{(*)0}\rangle=|1/2,1/2\rangle$, and $|B^{(*)-}\rangle=-|1/2,-1/2\rangle$. Here, we only investigate the isospin $I=0$ sectors, since the interaction potentails of the $I=1$ sectors are repulsive for these systems and cannot expect any bound states appeared. In addition, the three parameters in Table 3 are taken as $\lambda_{c}\approx 1/4$ Debastiani:2017ewu ; Wang:2022aga , $\lambda_{cc}\approx 1/9$ Marse-Valera:2022khy , and $\lambda_{b}\approx 1/10$ Dias:2019klk , which are the suppression factors of the exchanged heavy vector mesons, $D_{(s)}^{*}$, $J/\psi$, and $B_{(s)}^{*}$ with respect to the exchanged light vector mesons, respectively. The contributions from the exchange of heavier vector mesons, such as $B_{c}^{*}$ and $\Upsilon$, are neglected. It is significant to note that the dominant contributions come from exchanging the light vector mesons. More discussions can be found in Refs. Debastiani:2017ewu ; Roca:2024nsi .
@@
This equation is divergent and we can use the three-momentum cutoff approach Oset:1997it or the dimensional regularization method Oller:2000fj ; Jido:2003cb . The expression for the three-momentum cutoff approach is given by
@@
which has two free parameters, the regularization scale $\mu$ and the subtraction constant $a(\mu)$, and where $q_{cml}(s)$ is the three momentum of the particle in the center-of-mass frame
@@
which is applied for the purely molecular states. In the case of composite states, where the states contain not only the molecular components but also the other non-molecular components, this sum rule of Eq. (21) can be modified as Aceti:2012dd
@@ III Results
As discussed in the last section, there are two free parameters in our formalism, the regularization scale $\mu$ and the subtraction constant $a(\mu)$. Thus, we should first determine the values of the regularization scale $\mu$ and the subtraction constant $a(\mu)$ in the loop functions. At first, we take $\mu=q_{max}=800$ MeV Marse-Valera:2022khy ; Roca:2024nsi , and then use the three-momentum cutoff approach and the dimensional regularization method to match their values of the loop function at the threshold for a given channel to determine $a(\mu)$, as done in Ref. Oset:2001cn .
In the $\Omega_{ccc}$-like sector, we find four candidates for the three-charm molecular pentaquarks, as shown in Table 4, where the corresponding poles, couplings and compositeness for a given coupled channel are presented. For each pole, we use the $+$ to indicate that the corresponding coupled channel is closed, and the -- to specify the open channel accordingly, which is the decay channel. The largest coupling and compositeness are shown in bold, which indicates the most relevant channel in the most cases, in other word, the bound channel. Thus, the first two poles in Table 4, $5446.97$ MeV and $(5493.65-5.92i)$ MeV, are mainly coupled to the channels $D\Xi_{cc}$ and $D\Xi_{cc}^{*}$, respectively, indicating that they are the molecular states of these two channels with the binding energies of $42$ MeV and $49$ MeV, respectively, while the contributions of the channels $D_{s}\Omega_{cc}$ and $D_{s}\Omega_{cc}^{*}$ are also significant with large couplings and compositenesses. And the second state has a width of about $\Gamma=12$ MeV, which can decay into the $\eta\Omega_{ccc}$ channel. The last two poles, $5591.87$ MeV and $(5640.10-3.43i)$ MeV, are degenerate in $J^{P}=\frac{1}{2}^{-}$, $\frac{3}{2}^{-}$ and $J^{P}=\frac{1}{2}^{-}$, $\frac{3}{2}^{-}$, $\frac{5}{2}^{-}$, respectively, which are possibly qualified as the bound states of $D^{*}\Xi_{cc}$ and $D^{*}\Xi_{cc}^{*}$ with binding energies of $39$ MeV and $43$ MeV, respectively. One can see that the last state can decay into the $\omega\Omega_{ccc}$ channel with a width of about $\Gamma=7$ MeV. From the perspective of compositeness results, our predicted states are mostly pure molecular states with very small non-molecular components, since the sum of the compositenesses is nearly close to one.
@@
Furthermore, in order to understand the properties of the bound states, in Fig. 3 we also plot the changes in the masses and widths of the four states listed in Table 4 with respect to the free parameter $\mu$. Within a considerable range of the $\mu$ values, the poles of these four states are located below thethresholds of corresponding channel $D\Xi_{cc}$, $D\Xi_{cc}^{*}$, $D^{*}\Xi_{cc}$ and $D^{*}\Xi_{cc}^{*}$, respectively. As the value of $\mu$ increases, the masses decrease, while the widths first increase and then decrease. From the results of Fig. 3, one can see that these systems are stably bound for a reasonable range of the $\mu$ values. Thus, our prediction for these bound systems are valuable for future experiments.
@@
Note that there are some predictions from the other work. For the $\Omega_{ccc}$-like sector, a $D\Xi_{cc}$ bound state with the mass about $4310-4330$ MeV was predicted in Ref. Hofmann:2005sw , which was at least about $150$ MeV below the threshold. In Ref. Romanets:2012hm , a similar state of the mass at $4358.2$ MeV was found, which mainly coupled to the $D\Xi_{cc}$ channel. These two predictions are more bound than what we obtain for the $D\Xi_{cc}$ state with a mass $5446.97$ MeV. Using the one-boson-exchange model, Ref. Chen:2017jjn found two candidates for the molecules of $D\Xi_{cc}$ and $D^{*}\Xi_{cc}$ with the binding energies of $0.22$ MeV and $18.71$ MeV, respectively. Compared to their results, the states we obtained are more bound, compared with our binding energy about $40$ MeV. For the $\Omega_{bbb}$-like sector, a $\bar{B}\Xi_{bb}^{*}$ bound state at $15212.04$ MeV with quantum number $J^{P}=\frac{3}{2}^{-}$ was predicted in Ref. Dias:2019klk . It is below the threshold of the $\bar{B}\Xi_{bb}^{*}$ channel around $300$ MeV, which is much bigger than our binding energy of about $47$ MeV. The main reason is that we use different scheme to regularize the loop functions compared to theirs. In addition, there may be some differences between the particle masses used in Ref. Dias:2019klk and ours, which can be seen from the corresponding thresholds.

### v3

@@ Abstract
In this study, we explore the interactions between mesons and baryons in the open heavy sectors to identify potential triple-heavy molecular pentaquarks. We derive the meson-baryon interaction potentials using the vector meson exchange mechanism within the extended local hidden gauge formalism. The scattering amplitudes are computed by solving the coupled-channel Bethe-Salpeter equation, revealing several bound systems. By analyzing the poles of these amplitudes in the complex plane, we determine the masses and widths of these bound states. Additionally, we evaluate the couplings and compositeness of different channels within each bound system to assess their molecular characteristics. Our predictions include four $\Omega_{ccc}$-like states, four $\Omega_{bbb}$-like states, fourteen $\Omega_{bcc}$-like states, and ten $\Omega_{bbc}$-like states, which could be targets for future experimental investigations.
@@ I Introduction
Over the past two decades, advancements in high-energy physics have led to the discovery of numerous new hadronic states. Many of these states do not conform to the traditional classifications of mesons, which consist of a quark and an antiquark, or baryons, which are made up of three quarks or antiquarks Gell-Mann:1964ewy ; Zweig:1981pd . Instead, these states may be candidates for exotic forms such as glueballs, hybrids, and multiquark states, which have garnered significant interest from both experimentalists and theorists Liu:2013waa ; Hosaka:2016pey ; Chen:2016qju ; Richard:2016eis ; Lebed:2016hpi ; Olsen:2017bmm ; Guo:2017jvc ; Liu:2019zoy ; Brambilla:2019esw ; Meng:2022ozq ; Chen:2022asf ; Liu:2024uxn . Understanding these states is crucial for advancing our knowledge of the non-perturbative behavior of strong interactions.
Among these reported states, a notable group includes hidden-charm hadrons, such as the charmonium-like $XYZ$ states and $P_{c}/P_{cs}$ pentaquark-like states, which have sparked extensive discussion regarding molecular multiquark states. For example, the $X(3872)$, the first experimentally observed charmonium-like $XYZ$ state, was reported by the Belle collaboration in 2003 through the decay $B^{\pm}\rightarrow K^{\pm}\pi^{+}\pi^{-}J/\psi$ Belle:2003nnu . Its mass, close to the $D\bar{D}^{*}$ threshold, suggests it could be a $D\bar{D}^{*}$ molecular state, as proposed in several studies Swanson:2003tb ; Wong:2003xk ; Close:2003sg ; Voloshin:2003nt ; AlFiky:2005jd ; Liu:2008fh ; Thomas:2008ja ; Liu:2009qhy ; Lee:2009hy ; Gamermann:2009uq ; Braaten:2010mg ; Baru:2013rta ; Baru:2015nea ; Song:2023pdq .
In 2013, the BESIII and Belle collaborations reported the charge charmonium-like state $Z_{c}(3900)$, which is also near the $D\bar{D}^{*}$ threshold and can be interpreted as a molecular state BESIII:2013ris ; Belle:2013yex ; Guo:2013sya ; Wang:2013daa ; Aceti:2014uea . Later, in 2015, the LHCb collaboration discovered the states $P_{c}(4380)^{+}$ and $P_{c}(4450)^{+}$ in the $J/\psi p$ invariant mass spectrum from the decay $\Lambda_{b}^{0}\rightarrow J/\psi pK^{-}$ LHCb:2015yax ; LHCb:2016lve . These states, initially predicted as molecular pentaquarks Wu:2010jy ; Wu:2010vk ; Wang:2011rga ; Yang:2011wz ; Wu:2012md ; Xiao:2013yca ; Li:2014gra ; Karliner:2015ina , have since received considerable theoretical attention, as discussed in reviews Hosaka:2016pey ; Chen:2016qju ; Olsen:2017bmm ; Guo:2017jvc .
In 2019, the LHCb collaboration further refined their observations, revealing that the previously identified $P_{c}$ states were actually three distinct states: $P_{c}(4312)^{+}$, $P_{c}(4440)^{+}$, and $P_{c}(4457)^{+}$, which provided strong support for the molecular interpretation LHCb:2019kea ; Chen:2019asm ; Chen:2019bip ; Liu:2019tjn ; He:2019ify ; Xiao:2019mst ; Guo:2019kdc ; Xiao:2019aya .
More recently, in 2021, the LHCb collaboration reported the hidden-charm pentaquark structure $P_{cs}(4459)^{0}$ in the decay $\Xi_{b}^{-}\rightarrow J/\psi\Lambda K^{-}$ LHCb:2020jpq . Additionally, a new state, $P_{cs}(4338)^{0}$, containing a strange quark, was observed in the decay $B^{-}\rightarrow J/\psi\Lambda\bar{p}$ LHCb:2022ogu . These states are considered strong candidates for $\bar{D}^{*}\Xi_{c}$ and $\bar{D}\Xi_{c}$ molecular states, respectively Xiao:2021rgp ; Du:2021bgb ; Zhu:2021lhd ; Lu:2021irg ; Zou:2021sha ; Wang:2021itn ; Wu:2021caw ; Yan:2022wuz ; Meng:2022wgl ; Burns:2022uha ; Ortega:2022uyu . Thus, within the molecular framework and flavor symmetry, it is natural to assign the $P_{c}(4312)$ and $P_{c}(4440)$, $P_{c}(4457)$ as $\bar{D}\Sigma_{c}$ and $\bar{D}^{*}\Sigma_{c}$ states, respectively, and the $P_{cs}(4338)$ and $P_{cs}(4459)$ as $\bar{D}\Xi_{c}$ and $\bar{D}^{*}\Xi_{c}$ molecules, respectively, replacing the light quark with a strange quark, $q\to s$.
@@
From the molecular perspective, the hidden-charm pentaquark states $P_{c}$ and $P_{cs}$ can be seen as a straightforward substitution of $\bar{q}\rightarrow qq$ in the $XYZ$ states, as illustrated in Fig. 1. Similarly, by substituting $\bar{c}\rightarrow cc$ in the $XYZ$ states, we obtain triple-charm pentaquark states through the interaction between a charmed meson and a double-charm baryon, which is of significant interest.
In 2013, Ref. Guo:2013xga predicted the existence of exotic triple-heavy pentaquarks based on heavy antiquark–diquark symmetry, assuming that the $X(3872)$ and $Z_{b}(10610/10650)$ were molecular states. Subsequently, Refs. Chen:2017jjn ; Wang:2019aoc explored this further, predicting isoscalar triple-charm molecular-type pentaquark candidates such as $D^{(*)}\Xi_{cc}$, $D_{1}\Xi_{cc}$, and $D_{2}^{*}\Xi_{cc}$ using the one-boson-exchange model. These predictions offer valuable insights for future experiments.
Building on these predictions and the motivation from the findings of the $P_{c}$ and $P_{cs}$ states, we investigate triple-heavy pentaquark systems in the present work. Specifically, we consider pentaquark systems with quark contents $cccq\bar{q}$, $bbbq\bar{q}$, $bccq\bar{q}$, and $bbcq\bar{q}$, which we refer to as $\Omega_{ccc}$-like, $\Omega_{bbb}$-like, $\Omega_{bcc}$-like, and $\Omega_{bbc}$-like molecular states. For the meson-baryon interaction, we use a coupled-channel approach with transition potentials derived from the extended local hidden gauge formalism. Previous studies Wang:2022aga ; Wang:2023mdj ; Song:2024yli have predicted many molecular pentaquark states with double heavy quarks using this formalism.
Our paper is constructed as follows. In Sec. II, we introduce the $S$-wave interactions of the meson-baryon derived from the extended local hidden gauge formalism. Then, in Sec. III, we give the results of the triple-heavy pentaquark candidates by solving the coupled channel Bethe-Salpeter equation. Finally, a short summary is shown in Sec. IV.
@@ II Formalism
Note that Eqs. (3) and (4) are the matrices of $q\bar{q}$ written in terms of mesons and that the results emerging from Eqs. (1) and (2), can be obtained simply using the $q\bar{q}$ content of the mesons without invoking SU(5) symmetry as shown in Sakai:2017avl . Therefore, the vertices $VPP$ and $VVV$ can be obtained from the effective Lagrangians of Eqs. (1) and (2).
However, for the vertex of $VBB$, it is difficult to directly extend the corresponding baryon fields matrix from the flavor SU(3) notation to the one of SU(5) with the charm and beauty baryon fields. Therefore, instead of the evaluation from the effective Lagrangian, we use the scheme of Ref. Debastiani:2017ewu to calculate the $VBB$ vertex from the flavor and spin wave functions of baryons used in Capstick:1986ter ; Roberts:2007ni in which the heavy quarks are singled out and the symmetry of identical particles is imposed on the light quarks. The baryons involved in our work, along with their explicit flavor and spin wave functions, are listed in Table 2, where $\chi_{MS}$ indicates that the spin wave function is mixed symmetric, and $\chi_{MA}$ represents the mixed antisymmetric spin wave function. For the $J^{P}=\frac{1}{2}^{+}$ ground baryon states in the case of $S_{z}=+\frac{1}{2}$, the corresponding spin wave functions are given by
@@
Based on the local hidden gauge Lagrangians for the $VPP$ and $VVV$ vertices discussed above, and the extended approach for the $VBB$ vertex with the baryon wave functions, we obtain the general form of the tree level potentials after projecting to the $S$ wave
@@
Note that the coefficients of $C_{ij}$ in Table 3 are consistent with our isospin notation, where we use the phase convention $|D^{(*)+}\rangle=|1/2,1/2\rangle$, $|D^{(*)0}\rangle=-|1/2,-1/2\rangle$, $|\bar{B}^{(*)0}\rangle=|1/2,1/2\rangle$, and $|B^{(*)-}\rangle=-|1/2,-1/2\rangle$. Here, we only investigate the isospin $I=0$ sectors, since the interaction potentials of the $I=1$ sectors are repulsive for these systems and bound states do not appear. In addition, the three parameters in Table 3 are taken as $\lambda_{c}\approx 1/4$ Debastiani:2017ewu ; Wang:2022aga , $\lambda_{cc}\approx 1/9$ Marse-Valera:2022khy , and $\lambda_{b}\approx 1/10$ Dias:2019klk , which are the suppression factors of the exchanged heavy vector mesons, $D_{(s)}^{*}$, $J/\psi$, and $B_{(s)}^{*}$ with respect to the exchanged light vector mesons, respectively. The contributions from the exchange of heavier vector mesons, such as $B_{c}^{*}$ and $\Upsilon$, are neglected. It is significant to note that the dominant contributions come from exchanging the light vector mesons. More discussions can be found in Refs. Debastiani:2017ewu ; Roca:2024nsi .
@@
This magnitude is divergent and we can use the three-momentum cutoff approach Oset:1997it or the dimensional regularization method Oller:2000fj ; Jido:2003cb . The expression for the three-momentum cutoff approach is given by
@@
which also has only one free parameter the regularization scale $\mu$, and the subtraction constant $a(\mu)$ depends on the $\mu$ chosen, and where $q_{cml}(s)$ is the three momentum of the particle in the center-of-mass frame
@@
which is applied for the purely molecular states, where each of the terms on the left hand side of Eq. (21) gives the probability of each channel. In the case of composite states, where the states contain not only the molecular components but also the other non-molecular components, this sum rule of Eq. (21) can be modified as Aceti:2012dd
@@ III Results
As discussed in the last section, there are only one free parameter in our formalism, the regularization scale $\mu$, and the subtraction constant $a(\mu)$ depends on the $\mu$ chosen. Thus, we should first determine the values of the regularization scale $\mu$ and the subtraction constant $a(\mu)$ in the loop functions. At first, we take $\mu=q_{max}=800$ MeV Marse-Valera:2022khy ; Roca:2024nsi , and then use the three-momentum cutoff approach and the dimensional regularization method to match their values of the loop function at the threshold for a given channel to determine $a(\mu)$, as done in Ref. Oset:2001cn .
In the $\Omega_{ccc}$-like sector, we find four candidates for the three-charm molecular pentaquarks, as shown in Table 4, where the corresponding poles, couplings and compositeness for a given coupled channel are presented. For each pole, we use the $+$ to indicate that the corresponding coupled channel is closed, and the -- to specify the open channel accordingly, which is the decay channel. The largest coupling and compositeness are shown in bold, which indicates the most relevant channel in most cases, in other words, the most relevant bound channel. Thus, the first two poles in Table 4, $5446.97$ MeV and $(5493.65-5.92i)$ MeV, are mainly coupled to the channels $D\Xi_{cc}$ and $D\Xi_{cc}^{*}$, respectively, indicating that they are the molecular states of these two channels with the binding energies of $42$ MeV and $49$ MeV, respectively, while the contributions of the channels $D_{s}\Omega_{cc}$ and $D_{s}\Omega_{cc}^{*}$ are also significant with large couplings and compositenesses. And the second state has a width of about $\Gamma=12$ MeV, coming from the decay into the $\eta\Omega_{ccc}$ channel. The last two poles, $5591.87$ MeV and $(5640.10-3.43i)$ MeV, are degenerate in $J^{P}=\frac{1}{2}^{-}$, $\frac{3}{2}^{-}$ and $J^{P}=\frac{1}{2}^{-}$, $\frac{3}{2}^{-}$, $\frac{5}{2}^{-}$, respectively, which are possibly qualified as the bound states of $D^{*}\Xi_{cc}$ and $D^{*}\Xi_{cc}^{*}$ with binding energies of $39$ MeV and $43$ MeV, respectively. One can see that the last state can decay into the $\omega\Omega_{ccc}$ channel with a width of about $\Gamma=7$ MeV. From the perspective of compositeness results, our predicted states are mostly pure molecular states with very small non-molecular components, since the sum of the compositenesses is nearly close to one.
@@
Furthermore, in order to understand the properties of the bound states, in Fig. 3 we also plot the changes in the masses and widths of the four states listed in Table 4 with respect to the free parameter $\mu$. Within a considerable range of the $\mu$ values, the poles of these four states are located below the thethresholds of corresponding channel $D\Xi_{cc}$, $D\Xi_{cc}^{*}$, $D^{*}\Xi_{cc}$ and $D^{*}\Xi_{cc}^{*}$, respectively. As the value of $\mu$ increases, the masses decrease, while the widths first increase and then decrease. From the results of Fig. 3, one can see that these systems are stably bound for a reasonable range of the $\mu$ values. Thus, our prediction for these bound systems are valuable for future experiments.
@@
Note that there are some predictions from the other work. For the $\Omega_{ccc}$-like sector, a $D\Xi_{cc}$ bound state with the mass about $4310-4330$ MeV was predicted in Ref. Hofmann:2005sw , which was at least about $150$ MeV below the threshold. In Ref. Romanets:2012hm , a similar state with the mass at $4358.2$ MeV was found, which mainly coupled to the $D\Xi_{cc}$ channel. These two predictions are more bound than what we obtain for the $D\Xi_{cc}$ state with a mass $5446.97$ MeV. Using the one-boson-exchange model, Ref. Chen:2017jjn found two candidates for the molecules of $D\Xi_{cc}$ and $D^{*}\Xi_{cc}$ with the binding energies of $0.22$ MeV and $18.71$ MeV, respectively. Compared to their results, the states we obtained are more bound, compared with our binding energy about $40$ MeV. For the $\Omega_{bbb}$-like sector, a $\bar{B}\Xi_{bb}^{*}$ bound state at $15212.04$ MeV with quantum number $J^{P}=\frac{3}{2}^{-}$ was predicted in Ref. Dias:2019klk . It is below the threshold of the $\bar{B}\Xi_{bb}^{*}$ channel around $300$ MeV, which is much bigger than our binding energy of about $47$ MeV. The main reason is that we use different scheme to regularize the loop functions compared to theirs. In addition, there may be some differences between the particle masses used in Ref. Dias:2019klk and ours, which can be seen from the corresponding thresholds.